# An Adaptive Computational Intelligence Approach to Personalised Health and Immune Age Characterisation from Common Hematological Markers

### v5

@@ 4 Results
In Table 6, we provide examples of CCIS and its correspondent NCCIS values for a sample of real-life and artificial cases. The second column of the table shows the NIS, for purposes of comparison. It is worth noting that these examples were generated using the NHS values for normal healthy subjects, and based on the provisional assumption that the distance between the upper and normal ranges is twice the standard deviation. The examples were taken from [7] as they were intended as very preliminary tests. We are aware that their source clearly states they are meant only for teaching purposes. Here they are used for purposes of illustration only, and the development of the NIS and NCCIS is not dependent on them.
@@
One can observe how the NCCIS values meet the requirements we set in advance of their definition. We tested the NIS and NCCIS against real data from the National Health and Nutrition Examination Survey 2003–2016 (NHANES), provided in [28]. One measure of success was their ability to show how accurately the score discriminates between healthy subjects and patients suffering from various diagnosed illnesses.
@@
To this end, we examined thousands of cases from [28] and calculated the distribution, mean values, and standard deviations for each of the analytes (see Fig. 7. We used the means and standard deviations from the data to define the metrics for both the NIS and NCCIS. Finally, we calculated the NIS and NCCIS for thousands of cases in the NHANES database to find individual scores. The resulting values were grouped as belonging to healthy individuals or to a selected list of common diseases.
@@
Figure 4 illustrates the distribution of NIS values across various (self-declared) conditions for individuals in the NHANES database. The table 4 presents the number of data points used, along with the respective ANOVA results to test statistical significance. It must be noted that the self-reported conditions were not independently confirmed. Moreover, the survey participants did not distinguish between current or past diagnoses. This could be expected to introduce some noise, as some currently healthy people will be labelled as having a condition and some people with conditions will not have been diagnosed. We hope to improve accuracy by filtering the data or by adding other data sources in the future.
@@
Conversely, the NCCIS did not substantially improve our knowledge of how some diseases impact analyte counts. There may be underlying co-morbidities or cases of blood-immune disorders that could create anomalous deviations from the predictions. Regardless, the methods are shown to be valid for predicting the complex immune profiles/trends (variation over time) observed in diverse patients when considering socio-clinical metadata parameters such as age, and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to such deviant clinical cases in prospective studies. In Fig. 5, it can be seen that reported medical conditions produce a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We thus concluded that the NCCIS did enhance the knowledge already gleaned from the NIS and the CCIS.
@@ 4.1 Linearisation and Separability
For example, the neutrophil-to-lymphocyte ratio relation is a key determinant of severity for sepsis, and involves 2 analytes. The synthetic analyte that can be added is the ratio itself, thus replacing a non-linear rule that would eventually make the score’s description too convoluted to read with ease with a key marker as another analyte [20, 25].
@@
Figure 6: Distributions within ‘healthy’ reference values are heavily skewed in most cases, do not follow a normal or uniform distribution and often have long tails suggesting a personalised approach to reference values for abnormal results. The list of analyte labels is in the Supplementary Information.
@@
We observed that the self-reported health status of the patients did not accurately predict their clinically assessed blood analytes. This was expected because self-reports may have reflected illness at various points in time in a person’s life rather than ongoing conditions. In other words, with this experiment, we de-noised the data by disregarding self-reported status to test the score against cleansed data based exclusively on CBC results as a discriminant, according to medical guidance on normal population reference values. That is, the interpretation of the health status a person could have potentially received from a medical doctor based exclusively on their CBC.
When healthy and unhealthy patients were separated using a subset of four CBC/FBC analyte measures, namely, MCH, HCT, neutrophil, and lymphocyte counts, as shown in Figure 13, we obtained the second best $R^{2}$ value of 0.075, for both sexes. The $R^{2}$values were 0.050 and 0.123, for the male and female groups, respectively. We found that neutrophil count alone, as a single clinical parameter, had a near equal performance as the composite measure of these four analyte parameters, with an $R^{2}$ of 0.072 when including both sexes. The inclusion of other white blood cell (WBC) parameters, such as monocyte count, eosinophil, and basophil, did not outperform these correlates. We observed that different subsets of a CBC/FBC by health status separated the data. All reported measures were statistically significant, with a two-tailed P<0.0001<0.0001.
This also suggests that the mutlivariate biopsychosocial determinants of health should also be integrated for a top-bottom analysis of the disease predisposition and health dynamics of patients when assessing such self-reported measures. The irreducibility (multidimensionality) of these self-reported health measures may help explain why they are highly noisy and intersect with multiple other health factors, including multi-scale stressors such as the time of day, basic survival factors such as nutritional status/intake, sleep-related mood, psychological state, and social determinants of health, together with other health parameters.
@@ 4.2 Adaptive reference values
The main underlying assumption of this adaptive immune score definition is that, if a significant event (in the clinical sense) can be measured via an analyte, such an event will present itself as a value that breaks from the trend defined by the behaviour of the analyte over time. A second assumption is that such events can occur within the normal range as defined by a population reference range. A third assumption is that, while a trend becomes more incontrovertible the more data points there are, past values are less significant than proximate events. The next assumption is that the trend can be non-monotonous. The final assumption is that standard reference ranges can detect such events with acute sensitivity, while low specificity can thus be improved by narrowing the ranges but not by expanding them.
@@
The function described in the preceding paragraph represents the adaptive mean, from which the adaptive ranges are inferred. The adaptive ranges consist of two functions, the upper limit and the lower limit. The upper limit is defined as the adaptive mean plus the standard deviation of all the values that are above the mean, up to the point of time in question. Similarly, the lower limit is the adaptive mean minus the standard deviation of all the values that are below the adaptive mean up to the relevant point in time. Next, we smooth the transition from the population ranges to the adaptive ranges by means of a linear combination of both ranges such that the adaptive range becomes the dominant factor in the time function.
@@
Figure 7: Statistical inference without smart outlier detection. In contrast to the smart outlier detection technology, simple statistical regression towards the observed values results in skewed inferred ranges.
Figure 8: Smart outlier detection example. The machine learning technology can adapt the range of expected values for all analytes, showing stable trends, while also detecting abnormal values or trends. In the image, we can show the result of machine learning inference of the ranges of normal values for a patient while also detecting abnormal values.
@@ 5 Immune age estimation
As lifespan increases in many countries, there is a concomitant increase in age-related morbidities, particularly cardiovascular diseases, and increased susceptibility to seasonal infections, cancers, and neurodegenerative disorders. A unifying link between the higher rates of these disparate diseases observed in ageing populations is the progressive decline in immune functions.
@@
Here we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, gives insight into the immune age as compared to chronological age. Chronological age simply indicates the age of the individual, as determined by the length of time someone has lived or is living since their birth. This term should not be confused in the context of other literature, such as age estimates predicted by DNA methylation patterns from the Horvath algorithm.
@@ 5.1 Immune Score Over time
We hypothesised that as the human body ages the immune score will vary over time. In particular, we expected to see an upwards trajectory in direct relation to age.
To test the hypothesis, we selected two cohorts comprising all the data entries of individuals with ages ranging between 20 to 90 years, from the dataset, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex, we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
Figure 9: (a) The mean immune age as a function of age for NHANES data. The reference ranges were computed over the population aged 20 to 85 for males and females, respectively. Each of the data points represents the mean immune score for the cohort [age5,age+5][age-5,age+5]; for example, 50 represents the immune score for the population aged 45 to 55 years. (b)The distribution of immune age as a function of age. While a correlation trend is visible (especially for males), across sex the correlation is weak because healthy and non-healthy cohorts (both self-reported and by ground truth) are confounded.
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune score, as well as an increase in the 50th percentile and variance. The immune score decreases slightly for males between 20 and 25 year old. However, the trend in females is slowed down (compared to males) between ages 40 and 50 and reversed between ages 50 and 60. It remains to be investigated whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages where the slow-down and reversal in the age score trend for females is detected. It is known and has been reported that pregnancy and menopause produce a higher and more progressive increase in red cell counts, haematocrit levels and increased mean cell volume (MCV), and haemoglobin concentrations [21, 4, 9, 27, 1].
5.2 Immune age based on the immune score
@@
where $\alpha^{*}$ is the closest score in $IAC$ to $\alpha$. If there are two scores at the same distance, we choose the leftmost option. The resulting function is shown in Fig. 13. The curve shows some minor bumps which are the results of the discretisation used. We maintain that this distortion is not significant.
@@
As shown in Figure 14, prediction error rates are much lower in healthy NHANES data, compared to unhealthy groups, for both sexes and therefore can be used to estimate the chronological age of healthy patients. Prediction error rates for the immune score age were overestimated by approximately 7 years, for the healthy groups, in both sexes, allowing a correction (mean error rate of 7.33 $\pm$ 1.21 for healthy females, and 7.62 $\pm$ 1.17 for healthy males). Furthermore, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction. Pearson’s correlation statistics on the age error rates predicted, by the health status of the red blood cell count, as discussed before, were an R square value of 0.293 and 0.168, for men (P<< 0.0001) and females (P<<0.013). Age prediction error rates were high and not significant when self-reported health measures were used to separate health status groups, but were significant when taking abnormal CBC values only.
@@
BASO 0.0051 0.0166 0.0004
EOS 0.0012 2.00E-05 0.0049
HCT 0.0036 0.0055 0.0696
HGB 0.1179 0.1397 0.0635
LYMP 0.0429 0.4983 0.0928
MCH 0.0831 0.0001 0.1538
MCHC 0.0025 0.0043 0.0047
MCV 0.2661 0.337 0.2186
MONO 0.0121 0.0004 0.0073
NEUT 0.0084 0.0069 0.018
RBC 0.0268 1.00E-05 0.0009
PLT 0.0921 0.3334 0.1029
Table 7 shows the statistics of the predictors and separators of age for healthy and unhealthy populations. These plots show that common blood marker dynamics are predictive of age and health.
There are small but potentially meaningful differences between reference ranges. This could affect the classification of out-of-range counts when applying mismatched country-specific ranges. Using uniform, population-appropriate references is important for generalisable models but our main contribution is to learn each individual personal reference values hence making the reference discrepancies per country less relevant.
@@
Figure 11: Age Prediction Error Rates between actual patient (biological) age and Immune-score predicted age from NHANES dataset by RBC counts. The health status was determined by the best blood analyte correlate in this case, as previously discussed, the RBC counts, to separate healthy from unhealthy groups by sex. The results show that the immune-score predicted age has the lowest prediction error rate among healthy groups, for both sexes and overestimates age when the individuals are not healthy.

### v7

@@ 3 Results
Figure 3: Immune score space: The $x$ axis represents the deviation of all the analytes from the expected value, while the $y$ axis measures the total number of analytes. A negative standard deviation indicates that the value is lower than expected and a positive one indicates a higher than expected value. The cut-off seen on the negative $x$ axis shows the fact that real-life values for analytes have a strict lower bound but, technically, have no upper bounds. For instance, it is not possible to have a negative white blood cell count.
Figure 4: The behaviour of NCCIS over synthetically generated patients with differing values and a deviation from the norm with regard to the respective expected values. The $x$ axis measures the stated analyte, while the $y$ axis measures the total number of analytes. In contrast to the behaviour of the immune score (Fig. 3), NCCISS is a non-linear function that displays a step-wise behaviour by design, while remaining independent of the number of analytes.
Comparison Sample Size (n) p-value
Healthy vs Unhealthy 7007 vs 6074 P << 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P << 0.0001
Healthy vs Anaemia 7007 vs 980 P << 0.0001
Healthy vs HIV 7007 vs 62 P << 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P << 0.0001
Healthy vs Leukaemia 7007 vs 18 P << 0.0001
Healthy vs Blood Cancer 7007 vs 8 P << 0.0001
Table 4: Differences in NIS mean values. All comparisons were performed using a one-way ANOVA.
Figure 5: Distribution of the Normalised Immune Score (NIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2014 database. The Healthy cohort is defined as those individuals with all analytes within the normal ranges as defined by the NHS (1); others comprise the Unhealthy set. The Self rep. healthy cohort comprises individuals who self-reported health of 3 (”Good”) or less. The number of individuals in each cohort is shown in table 4.
Comparison Sample Size (n) p-value
Healthy vs Unhealthy 7007 vs 6074 P << 0.0001
Self rep. Healthy vs Unhealthy 11976 vs 6074 P << 0.0001
Healthy vs Anaemia 7007 vs 980 P << 0.0001
Healthy vs HIV 7007 vs 62 P << 0.0001
Healthy vs Hodgkin’s Disease 7007 vs 37 P << 0.0001
Healthy vs Leukaemia 7007 vs 18 P << 0.0001
Healthy vs Blood Cancer 7007 vs 8 P << 0.0001
Table 5: Differences in NCCIS mean values. All comparisons were performed using a one-way ANOVA. Effect of sample size was analysed in Table 6 for statistical significance and practical applicability.
Parameter Cohen’s d p-value
IMMUNE AGE (by Health) 2.28 P << 0.0001
SCORE 1.36 P << 0.0001
IMMUNE AGE (by Sex) 0.11 P << 0.0001
Table 6: Cohen’s d effect sizes of combined blood parameters in the NIS immune Score, and Immune Age comparing healthy and unhealthy groups. Effect sizes highlight significant differences between the two groups. P-values were computed using an independent Welch’s t-test, which compares the means of two groups. Results are statistically significant, suggesting robust differences. The Immune Score (SCORE) indicates a discriminatory distinction between the two groups for practical purposes with SCORE, the immune score, and immune age having a statistical power above 1.00 (alpha 0.05) suggesting a low risk of false discovery. The Immune Age by Health status shows a strong Cohen’s d (2.28), indicating it is a highly robust measure for differentiating health status and predicting blood parameter-based health risks with precision. In contrast, the weaker effect size by Sex (0.11) suggests that sex is unlikely to be a confounding factor, reinforcing the accuracy of the Immune Age prediction based on health status.
Figure 6: Distribution of the Normalised Colour-Coded Immune Score (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000-2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (1) and Unhealthy otherwise. Self rep. healthy comprises individuals who self-reported health of 3 (“Good”) or less. The number of individuals in each cohort is shown in table 5. One explanation for the bimodality of some of these plots is the nature of the source data as self-reported, where respondents may or may not have been ill at the time of giving their answers, or from the application of the separating cutoff value from NIS. In either case, both score-separating capabilities are highly statistically significant. The separation of groups of self-reported data suggests that they could be separated from those that were actually not healthy at the time of testing. In contrast, those self-reported not healthy were indeed clinically not healthy according to their results and the score.
@@
In Table 7, we provide examples of CCIS and its correspondent NCCIS values for a sample of real-life and artificial cases. The second column of the table shows the NIS, for purposes of comparison. It is worth noting that these examples were generated using the NHS values for normal healthy subjects, and based on the provisional assumption that the distance between the upper and normal ranges is twice the standard deviation. The examples were taken from [9] as they were intended as very preliminary tests. We are aware that their source clearly states they are meant only for teaching purposes. Here they are used for purposes of illustration only, and the development of the NIS and NCCIS is not dependent on them. In Table 6, we provide Cohen’s d values suggesting practical applicability of the score, making it clinically relevant for triaging purposes even at the level of individual classification and prediction.
@@
One can observe how the NCCIS values meet the requirements we set in advance of their definition. We tested the NIS and NCCIS against real data from the National Health and Nutrition Examination Survey 2003-2016 (NHANES), provided in [32]. One measure of success was the ability to show how accurately the score discriminates between healthy subjects and patients suffering from various diagnosed illnesses.
@@
To this end, we examined thousands of cases from [32] and calculated the distribution, mean values, and standard deviations for each of the analytes (see Fig. 7. We used the means and standard deviations from the data to define the metrics for both the NIS and NCCIS. Finally, we calculated the NIS and NCCIS for thousands of cases in the NHANES database to find individual scores. The resulting values were grouped as belonging to healthy individuals or to a selected list of common diseases. Furthermore, Figure 8 demonstrates how machine learning predictions can dynamically adjust the expected range of analyte values, detecting stable trends while identifying outliers, thus enhancing precision in monitoring patient health.
@@
Figure 5 illustrates the distribution of NIS values across various (self-declared) conditions for individuals in the NHANES database. The table 4 presents the number of data points used, along with the respective ANOVA results to test statistical significance. In principle, smaller sample sizes in one group could lead to unstable variability in p-values. However, the consistent statistical significance across all comparisons and the large effect sizes shown in Table 6 indicate that the results are robust. Furthermore, Cohen’s d is based on the entire dataset, using group means, variances, and sample sizes to compute the pooled standard deviation, thereby accounting for sample sizes in the reported effect sizes. The large effect size of the immune age indicates it significantly distinguishes healthy from unhealthy individuals, reflecting its statistical and practical significance in health status discrimination and prediction. It must be noted that the self-reported conditions were not independently confirmed. Moreover, the survey participants did not distinguish between current or past diagnoses. This could be expected to introduce some noise, as some currently healthy people will be labelled as having a condition and some people with conditions will not have been diagnosed. We hope to improve accuracy by filtering the data or by adding other data sources in the future.
@@
The method is shown to be valid for predicting immune profiles/trends (variation over time) observed in diverse patients when considering socio-clinical metadata parameters such as age and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to deviant clinical cases in prospective studies. In Fig. 6, it can be seen that reported medical conditions produced a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We thus concluded that the NCCIS did enhance the insights already gleaned from the NIS and the CCIS.
@@
Test Statistic p-value
Kruskal-Wallis Test (Across all three scores) H-statistic = 10.49 0.0053
Pearson Correlation (NIS vs NCCIS) Coefficient = 0.89 << 0.0001
Wilcoxon Test (NIS vs NCCIS) N/A 0.018
Table 8: Statistical test results for the three scores: NIS, CCIS and NCCIS.
Statistical tests assessed the relationship among the three individualised scores, NIS, NCCIS, and CCIS, across various health conditions. Table 8 presents significant statistical associations between these scores. The Kruskal-Wallis test, a non-parametric alternative to ANOVA was performed due to the ordinal nature of the CCIS score. The significant p-value (0.0053) indicates that there are statistically significant differences in the distribution of NIS, NCCIS, and CCIS scores across the conditions. A Pearson correlation assessed the linear relationship between NIS and NCCIS. The strong positive correlation coefficient (0.89) with a highly significant p-value (¡ 0.0001) determined from a two-tailed independent samples t-test suggests a robust linear association between these two scores. Lastly, a Wilcoxon Signed-Rank test compared paired differences between NIS and NCCIS. The significant result (p = 0.018) indicates that there are significant differences between these paired scores.
@@ 3.1 Linearisation and Separability
For example, the neutrophil-to-lymphocyte ratio relation is a key determinant of severity for sepsis, and involves 2 analytes. The synthetic analyte that can be added is the ratio itself, thus replacing a non-linear rule that would eventually make the score’s description too convoluted to read with ease with
@@
We observed that the self-reported health status of the patients did not accurately predict their clinically assessed blood analytes. This was expected because self-reports may have reflected illness at various points in time in a person’s life rather than ongoing conditions. In other words, with this experiment, we de-noised the data by disregarding self-reported status to test the score against cleansed data based exclusively on CBC results as a discriminant, according to medical guidance on normal population reference values, that is, the interpretation of their health status a person could have potentially received from a medical doctor based exclusively on their CBC.
When healthy and unhealthy patients were separated using a subset of four CBC/FBC analyte measures, namely, MCH, HCT, neutrophil, and lymphocyte counts, as shown in Figure 13, we obtained the second best $R^{2}$ value of 0.075, for both sexes. The $R^{2}$ values were 0.050 and 0.123, for the male and female groups, respectively. We found that neutrophil count alone, as a single clinical parameter, delivered a performance nearly equal to that of the composite measure of these four analyte parameters, with an $R^{2}$ of 0.072 when including both sexes. The inclusion of other white blood cell (WBC) parameters, such as monocyte count, eosinophil, and basophil, did not outperform these correlates. We observed that different subsets of a CBC/FBC by health status separated the data. All reported measures were statistically significant, with a two-way ANOVA P<0.0001<0.0001.
This further suggests that the mutlivariate biopsychosocial determinants of health should also be integrated for a top-bottom analysis of the disease predisposition and health dynamics of patients when assessing such self-reported measures. The irreducibility (multidimensionality) of these self-reported health measures may help explain why they are highly noisy and intersect with multiple other health factors, including multi-scale stressors such as the time of day, basic survival factors such as nutritional status/intake, sleep-related mood, psychological state, and social determinants of health, together with other health parameters.
@@ 3.2 Adaptive Reference Values
The main underlying assumption of this adaptive immune score definition is that if a significant event (in the clinical sense) can be measured via an analyte, such an event will present itself as a value that breaks from the trend defined by the behaviour of the analyte over time. A second assumption is that such events can occur within the normal range as defined by a population reference range. A third assumption is that while a trend becomes more incontrovertible the more data points there are, past values are less significant than proximate events. The next assumption is that the trend can be non-monotonous. The final assumption is that standard reference ranges can detect such events with acute sensitivity, while low specificity can thus be improved upon by narrowing the ranges but not by expanding them.
@@
The function described in the preceding paragraph represents the adaptive mean, from which the adaptive ranges are inferred. The adaptive ranges consist of two functions, the upper limit and the lower limit. The upper limit is defined as the adaptive mean plus the standard deviation of all the values that are above the mean, up to the point of time in question. Similarly, the lower limit is the adaptive mean minus the standard deviation of all the values that are below the adaptive mean up to the relevant point in time. Next, we smooth the transition from the population ranges to the adaptive ranges by means of a linear combination of both ranges, such that the adaptive range becomes the dominant factor in the time function.
@@
Figure 7: Personalised reference values: Statistical inference without smart outlier detection. In contrast to the smart outlier detection technology, simple statistical regression towards the observed values results in skewed inferred ranges.
Figure 8: Smart outlier detection example. The machine learning technology can adapt the range of expected values for all analytes, showing stable trends while also detecting abnormal values or trends. In the image, we can show the result of machine learning inference of the ranges of normal values for a patient while also detecting abnormal values.
@@ 4 Immune Age Estimation
As lifespan increases in many countries, there is a concomitant increase in age-related morbidities, particularly cardiovascular disease, and increased susceptibility to seasonal infections, cancers, and neurodegenerative disorders. A unifying link between the higher rates of these disparate diseases observed in ageing populations is the progressive decline in immune function.
@@
Here we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, affords insight into the immune age as compared to chronological age. Chronological age simply indicates the age of the individual, as determined by the length of time someone has lived or has been living since their birth. This term should not be confused in the context of other literature, such as age estimates predicted by DNA methylation patterns from the Horvath algorithm.
@@ 4.1 Immune Score Over time
We hypothesised that as the human body aged the immune score would vary over time. In particular, we expected to see an upwards trajectory in direct relation to age.
To test the hypothesis, we selected from the dataset two cohorts comprising all the data entries of individuals with ages ranging between 20 and 90 years, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex, we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
Figure 9: (a) The mean immune age as a function of age for NHANES data. The reference ranges were computed over the population aged 20 to 85 for males and females, respectively. Each of the data points represents the mean immune score for the cohort [age5,age+5][age-5,age+5]; for example, 50 represents the immune score for the population aged 45 to 55 years. (b)The distribution of immune age as a function of age. While a correlation trend is visible (especially for males), across sex the correlation is weak because healthy and non-healthy cohorts are confounded at the source given that both and the patient’s actual current unhealthy states are in the same category.
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune score, as well as an increase in the 50th percentile and variance. The immune score decreases slightly for males between 20 and 25 year old. However, the trend in females is slowed down (compared to males) between ages 40 and 50 and reversed between ages 50 and 60. It remains to be investigated whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages where the slow-down and reversal in the age score trend for females is detected. It is known and has been reported that pregnancy and menopause produce a higher and more progressive increase in red cell counts, haematocrit levels and increased mean cell volume (MCV) as well as haemoglobin concentrations [25, 6, 11, 31, 2].
4.2 Discounting Confounding Factors in Immune Age
@@
where $\alpha^{*}$ is the closest score in $IAC$ to $\alpha$. If there are two scores at the same distance, we choose the leftmost option. The resulting function is shown in Fig.  10. The curve shows some minor bumps resulting from the discretisation used. We maintain that this distortion is not significant.
@@
As shown in Figure  11, prediction error rates are much lower in healthy NHANES data, compared to unhealthy groups, for both sexes, and therefore can be used to estimate the chronological age of healthy patients. Prediction error rates for the immune score age were overestimated by approximately 7 years, for the healthy groups of both sexes, allowing a correction (mean error rate of 7.33 $\pm$ 1.21 for healthy females, and 7.62 $\pm$ 1.17 for healthy males). Furthermore, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction, determined by two-tailed, two-sample t-tests. Pearson’s correlation statistics on the age error rates predicted by the health status of the red blood cell count as discussed before, were an R square value of 0.293 and 0.168, for men (P<< 0.0001) and women (P<<0.013) respectively. Age prediction error rates were high and not significant when self-reported health measures were used to separate health status groups, but were significant when taking abnormal CBC values only.
@@
BASO 0.01 0.02 0.00
EOS 0.00 0.00 0.00
HCT 0.00 0.01 0.07
HGB 0.12 0.14 0.06
LYMP 0.04 0.50 0.09
MCH 0.08 0.00 0.15
MCHC 0.00 0.00 0.00
MCV 0.27 0.34 0.22
MONO 0.01 0.00 0.01
NEUT 0.01 0.01 0.02
RBC 0.03 0.00 0.00
PLT 0.09 0.33 0.10
Table  LABEL:analytes shows the statistics of the predictors and separators of age for healthy and unhealthy populations. These plots show that common blood marker dynamics are predictive of age and health.
There are small but potentially meaningful differences between reference ranges. This could affect the classification of out-of-range counts when applying mismatched country-specific ranges. Using uniform, population-appropriate references is important for generalisable models, but our main contribution is to learn each individual’s personal reference values, thereby making the reference discrepancies per country less relevant. The best single analyte features for predicting health status and age are summarized in Table 10. Notably, MCV, LYMP, PLT, and HGB emerge as strong discriminators between healthy and unhealthy classes, each displaying distinct patterns across categories.
@@
Figure 11: Age Prediction Error Rates between actual patient (biological) age and Immune-score predicted age from NHANES dataset by RBC counts. The standard error analysis shown includes the prediction error rates in years (as box plots) across different health and sex groups, computed by assessing the variability in age prediction errors using interquartile ranges (IQR). The health status was determined by the best blood analyte correlate in this case as previously discussed– the RBC counts– to separate healthy from unhealthy groups by sex.
Fig. 11 highlights the fact that despite inter-group error rate variability, ML and computational methods can accurately differentiate between healthy and unhealthy groups, as evidenced by the well-separated box plots (medians and IQRs). The results show that the immune-score predicted age has the lowest prediction error rate among healthy groups, for both sexes, and overestimates age when the individuals are not healthy. Outliers beyond the 1.5 times IQR were not imputed to the nearest whisker point to showcase the variability and distribution of patient data.
4.3 Machine Learning ++ guiding score
Feature Model Accuracy MSE Sensitivity Specificity
LYMP RandomForest 1 0 1 1
LYMP KNN 1 0 1 1
NEUT RandomForest 1 0 1 1
NEUT KNN 1 0 1 1
HCT RandomForest 1 0 1 1
HCT KNN 1 0 1 1
MCH RandomForest 1 0 1 1
MCH KNN 1 0 1 1
Table 11: ML Performance metrics for Different Blood Analyte features and models (test size =0.2, with 10-fold cross-validation)
Model Accuracy MSE Sensitivity Specificity
RandomForest 1.00 0.00 1.00 1.00
SVM 1.00 0.00 0.96 1.00
KNN 1.00 0.00 0.99 1.00
Table 12: ML Performance metrics for different models using Blood Analytes
Model Accuracy MSE Sensitivity Specificity
RandomForest 0.98 0.02 0.15 0.99
SVM 0.99 0.01 0.06 1.00
KNN 0.99 0.01 0.08 1.00
ANN 0.99 0.01 0.04 1.00
Transformer-ANN 0.99 0.01 0.05 1.00
Table 13: ML Performance metrics for NIS Immune Score using Different models
Feature Model AUC
NIS Transformer-ANN 0.91
NIS ANN 0.91
NIS KNN 0.68
NIS RandomForest 0.70
NIS SVM 0.59
Table 14: AUC scores for each model under NIS Score.
Figure 12: ROC Curves for NIS Immune Score. The use of more complex ML algorithms like a feed-forward neural network (ANN) and a transformer-based attentional neural network shows an AUC above 0.9, enhancing the predictive performance of NIS immune scores. The higher performance of ANN and transformer models coupled with NIS scores suggests the ability of ANN systems to capture complex, non-linear relationships in the NIS scores, leading to enhanced predictive accuracy with computational intelligence.
As shown in Table  13 for the different AI models, the NIS immune score, a single quantitative measure (scalar value), also showed significant classification performance, particularly when coupled with computational AI-driven predictive power using ANN networks. While other basic ML classifiers did not perform as well, the ANN paired with the NIS immune score achieved an optimal predictive accuracy of 0.987, a high specificity of 0.999, and an AUC score of 0.91, comparable to the individual analytes themselves. However, the sensitivity was poor compared to the analytes. A low sensitivity in AI classification implies it is not able to correctly classify true positives. In contrast, the basic machine learning classifiers, such as RF, SVM, and KNN, did not perform as well as the ANN models with the NIS immune score, achieving lower AUC scores with 10-fold cross-validation, despite their high classification accuracy. While the feedforward ANN (MLP) detects local features through its dense layers, the Transformer-ANN captures long-range dependencies using self-attention mechanisms.
Table 12 shows the performance of all blood analytes collectively, and they perform sub-optimally vis-a-vis these four blood features alone. The score introduced captures in a single numerical value the power of this and other analytes that may be of interest for other patient profiles, also giving the clinician the opportunity and option to add or remove analytes or assign different weights to different analytes.Table 14 shows that both Transformer-ANN and ANN models demonstrate superior performance with an AUC of 0.91 for predicting outcomes based on the NIS score, significantly outperforming traditional models like KNN, RandomForest, and SVM.
Figure 12 illustrates these findings via ROC curves for the NIS immune score across different machine-learning models. The high AUC values achieved by the ANN models highlight their enhanced ability to capture complex, non-linear relationships in the NIS data. Therefore, they can significantly enhance predictive modeling and accuracy, suggesting their potential for improving diagnostic precision using immune scores.